KjSaif (kjsaif):

what could 1 +1 equal if it wasn't 2

OpenStudy (interstellar):

idk ;-;

OpenStudy (anonymous):

1+1 does equal 2 ?_?

OpenStudy (jadedry):

3 , due to rounding error. ; D

KjSaif (kjsaif):

i know but what if it didn't

OpenStudy (gottennis121):

Please close if you do not have a 13+ question. Thanks!

OpenStudy (jadedry):

alternatively, 1?

OpenStudy (anonymous):

11 lol

OpenStudy (anonymous):

see put 1 and a 1 together and you got 11 XD

OpenStudy (anonymous):

then maybe it could be like 1+ -1, 1+ 1. etc, but i believe 1+1 will always equal 2

OpenStudy (anonymous):

11

OpenStudy (pawanyadav):

It's 2(1)

KjSaif (kjsaif):

6? one + one lol

OpenStudy (pawanyadav):

That two times of one here we are taking 1 as variable BTW its a constant.

OpenStudy (anonymous):

dude gimme my medal you know you were looking for 11 ('-')/)

KjSaif (kjsaif):

lol

OpenStudy (sparrow2):

1+1=10 :)

KjSaif (kjsaif):

ha

OpenStudy (pawanyadav):

No, that's too different

KjSaif (kjsaif):

alright 3rd place is.............

KjSaif (kjsaif):

@Jadedry !

KjSaif (kjsaif):

2nd place is.......................

OpenStudy (jadedry):

third place, haha, ty

OpenStudy (anonymous):

i gave 3rd a medal tho

KjSaif (kjsaif):

@Pawanyadav because he was the first to make 1+1 real mathmatical XD

KjSaif (kjsaif):

and the winner is..............

KjSaif (kjsaif):

@Migoz300 for making the dumbest answer

OpenStudy (anonymous):

lol 2+2=

OpenStudy (anonymous):

triangle :D

OpenStudy (gottennis121):

lol

OpenStudy (pawanyadav):

2+2 =2(medals of Migos300)

OpenStudy (anonymous):

wait my bad 7+7= troangle 2+2= fish ;-;

RhondaSommer (rhondasommer):

IT IS SQUARE

OpenStudy (anonymous):

*triangle dang type pad ('-')/)